# Joint Channel and CFO Estimation From Beam-Swept Synchronization Signal Under Strong Inter-Cell Interference

### v1

@@ Abstract
Index Terms— CFO estimation, strong interference, beam-swept SS, multi-transmitter systems.
1 Introduction
@@ 2 System Model
Fig. 1: Preamble Structure.
2.1 Preamble Structure
@@
2.2 Transmission Model
For the $k$th BS, denote $\tau_{k}$ and $\omega_{k}$ as the group timing and frequency offsets respectively, and denote $\alpha_{k}^{(p)}$ denote the complex channel gain in the $i$th frame. Then, the received signal at time $m$ can be expressed as
@@ 2.2 Transmission Model
In this work, the initial tasks of active BS detection and delay estimation are assumed to be complete, consistent with established generalized likelihood ratio test (GLRT) methods [17]. The scope is therefore narrowed to estimating the remaining parameters for the active set $\mathcal{K}$: CFOs $\bm{\omega}\triangleq[\omega_{k}]_{k\in\mathcal{K}}$, scaling factors $\bm{\mu}\triangleq[\mu_{k}]_{k\in\mathcal{K}}$, and per-frame channel gains $\bm{\alpha}\triangleq[\alpha_{k}^{(p)}]_{k\in\mathcal{K},\,p\in\mathcal{P}}$.
$\underset{\bm{\omega},\bm{\alpha},\bm{\tau},\bm{\mu}}{\text{maximize}}\quad\mathbb{P}\left(\mathbf{y}|\bm{\omega},\bm{\alpha},\bm{\mu}\right).$
@@
3 Cross-Preamble Estimation
@@ 3.1 Channel Estimator
$\mathbf{y}^{(p)}=\underbrace{e^{j\omega(p-1)N_{\text{f}}}\bm{\Omega}\left(\bm{\omega}\right)\varodot\mathbf{C}\left(\bm{\tau}\right)}_{\bm{A}^{(p)}}\bm{\alpha}^{\left(p\right)}+\mathbf{v}^{(p)}$
where $\mathbf{C}\left(\bm{\tau}\right)=[\mathbf{0}_{\tau\times 1};\mathbf{c}_{k};\mathbf{0}]_{k\in\mathcal{K}}\in\mathbb{C}^{N_{\text{f}}\times K}$, $\bm{\Omega}\left(\bm{\omega}\right)=[\Omega_{k}(\omega_{k})]_{k\in\mathcal{K}}\in\mathbb{C}^{N_{\text{f}}\times K}$, $\Omega_{k}(\omega_{k})=[1;e^{j\omega_{k}};\cdots;e^{j\omega_{k}(N_{\text{f}}-1)}]$, $\bm{\alpha}^{(p)}=[\alpha_{k}^{(p)}]_{k\in\mathcal{K}}\in\mathbb{C}^{K\times 1}$, and $\mathbf{v}^{(p)}=[v[m]]_{m\in\{0,\cdots,N_{\text{f}}-1\}}\in\mathbb{C}^{N_{\text{f}}\times 1}$.
@@
$\mathbf{y}=\underbrace{\text{blkdiag}\left(\bm{A}^{(0)},\cdots,\bm{A}^{(P-1)}\right)}_{\bm{A}(\bm{\tau},\bm{\omega})}\bm{\alpha}+\bm{v}.$
Then, according to the LLS solution, the channel gain $\bm{\alpha}$ can be estimated by
$\hat{\bm{\alpha}}=\bm{A}(\bm{\tau},\bm{\omega})^{\dagger}\mathbf{y}.$ (5)
@@
$\bm{A}_{0}\bm{\alpha}+\underbrace{\left[\bm{A}^{(0)}\text{diag}\left(\bm{\alpha}^{\left(0\right)}\right);\cdots;\bm{A}^{(P-1)}\text{diag}\left(\bm{\alpha}^{\left(P-1\right)}\right)\right]}_{\bm{B}}\bm{\mu}+\bm{v}.$
where $\bm{A}_{i}=\text{blkdiag}(\bm{A}_{i}^{(0)},\cdots,\bm{A}_{i}^{(P-1)})$, $\bm{A}_{i}^{(p)}=e^{j\omega(p-1)N_{\text{f}}}$ $\bm{\Omega}(\bm{\omega})\varodot\mathbf{C}_{i}\left(\bm{\tau}\right)$, $\mathbf{C}_{i}\left(\bm{\tau}\right)=[\mathbf{0}_{(\tau+\tau_{\text{c}}i)\times 1};\mathbf{c}_{k,i};\mathbf{0}]_{k\in\mathcal{K}}$, and $\tau_{\text{c}}=N+\tau_{0}$. Then, according to the LLS solution, the scaling factor $\bm{\mu}$ can be estimated by
$\hat{\bm{\mu}}=\bm{B}^{\dagger}\left(\mathbf{y}-\bm{A}_{0}\right).$ (6)
3.2 Separate Analysis and Separate CFO Estimator
@@
$r_{y,k}^{p,i}\left[\tau\right]\triangleq\frac{1}{N}\sum_{m=\tau}^{\tau+N-1}y\left[m+pN_{\text{f}}+i\tau_{\text{c}}\right]c_{k,i}^{\text{H}}\left[m-\tau\right]$
@@
$r_{y,k}^{p,i}\left[\tau_{k}\right]\sim\alpha_{k}^{(p)}r_{k}\left(\omega_{k}\right)\text{$\left(\mu_{k}e^{-j\omega_{k}\tau_{\text{c}}}\right)$}^{i}+\mathcal{CN}\left(0,\sigma_{k,p,i}^{2}\right)$
@@
3.3 Cross-Preamble CFO Estimator
Let the set of correlation variables associated with CFO $\omega_{k}$ as $\mathbf{r}_{y,k}=[r_{y,k}^{p,i}\left[\tau_{k}\right]]_{p\in\mathcal{P},i\in\{0,1\}}$, and define the cross-preamble likelihood as $\mathbb{P}(\mathbf{r}_{y,k}|\bm{\omega})$. Since preambles from different frames are non-overlapping, the correlation outputs are conditionally independent given $\omega_{k}$, yielding
$\mathbb{P}\left(\mathbf{r}_{y,k}|\bm{\omega}\right)=\prod_{p\in\mathcal{P}}\prod_{i\in\{0,1\}}\mathbb{P}\left(r_{y,k}^{p,i}\left[\tau_{k}\right]|\omega_{k}\right).$
@@ 3.3 Cross-Preamble CFO Estimator
$\hat{\omega}_{k}^{\text{cross}}\triangleq\arg\max_{\omega_{k}}\mathbb{P}\left(\mathbf{r}_{y,k}|\bm{\omega}\right)=\frac{1}{\tau_{c}}\angle\left\{\psi_{1}\psi_{0}^{\text{H}}\right\}$ (9)
@@
0:Initialization: $\hat{\bm{\omega}}\leftarrow\mathbf{0}$, $\hat{\bm{\alpha}}\leftarrow\mathbf{0}$, $\hat{\bm{\mu}}\leftarrow\mathbf{1}$, $\tilde{\mathbf{y}}\leftarrow\mathbf{y}$
1: Construct $\mathbf{A}$ based on $\hat{\bm{\omega}}$ and $\hat{\bm{\mu}}$ and update $\hat{\bm{\alpha}}$ based on (5); Then, Construct $\mathbf{A}_{0}$ and $\mathbf{B}$ based on $\hat{\bm{\omega}}$ and $\hat{\bm{\alpha}}$ and update $\hat{\bm{\mu}}$ based on (6).
2: For each BS $k$, perform SIC and compensation: $\tilde{\mathbf{y}}_{k}\leftarrow(\mathbf{y}-\bm{A}(\hat{\bm{\tau}}_{-k},\hat{\bm{\omega}}_{-k})\hat{\bm{\alpha}}_{-k})\varodot\Omega_{k}(\omega_{k})$. Then, estimate the remaining CFO $\delta_{k}$ based on the cross-preamble estimator (9), and update $\hat{\omega}_{k}\leftarrow\hat{\omega}_{k}+\delta_{k}$.
@@
3.4 Joint CFO and Channel Estimation Algorithm
@@
$\tilde{\mathbf{y}}_{k}\;=\;\Bigl(\mathbf{y}-\bm{A}\bigl(\hat{\bm{\tau}}_{-k},\hat{\bm{\omega}}_{-k}\bigr)\hat{\bm{\alpha}}_{-k}\Bigr)\odot\bm{\Omega}_{k}(\hat{\omega}_{k})$ (13)
@@
4 Simulation Results
@@
Fig. 2 (a) illustrates the mean absolute error (MAE) over SINR. As expected, error decreases with SINR. The ranking is consistent with the length of the used reference sequence: cross-preamble < separate < CP-based. Cross-correlation outperforms auto-correlation because the separation $\tau_{\text{c}}$ between the two training sequences provides a larger phase baseline. The proposed cross-preamble estimator further improves accuracy by pooling information across frames under the shared-CFO constraint. Notably, the cross-preamble estimator at SINR $30\text{ dB}$ still outperforms the weighted-average baseline operating at SINR $5\text{ dB}$ confirming robustness in low-SINR regimes. Consequently, improving CFO accuracy reduces variance in channel estimation; the error (normalized mean squared error (NMSE)) versus SINR follows the same trend as in Fig. 2(b).
Fig. 2: CFO $\omega$ estimation MAE and channel $\alpha$ estimation NMSE over SINR, obtained from 200 Monte-Carlo simulations, with $K=12$ BSs, and $P=12$ preambles.
Fig. 3: Field test: Radio maps for two beams transmitted by a ground BS; the BS location is indicated by a red triangle.
Fig. 3 shows the radio-map estimation for a low-altitude setting. The two clearly resolved (pointing in different directions) beams corroborate the effectiveness of the proposed framework in practice.
5 Conclusion

### v2

@@ Abstract
I Introduction
@@ II System Model
Figure 1: Preamble Structure.
II-A Preamble Structure
@@
II-B Transmission Model
For the $k$th BS, denote $\tau_{k}$ and $\omega_{k}$ as the group timing and frequency offsets respectively, and denote $\alpha_{k}^{(p)}$ denote the complex channel gain in the $i$th frame. Then, the received signal at time $m$ can be expressed as [^1] The proposed algorithm can be naturally extended to frequency-selective fading by estimating the multipath gains and delays. However, this work focuses on the single-path (frequency-flat) case, which is sufficient for the low-altitude scenarios considered herein.
@@ II-B Transmission Model
In this work, the initial tasks of active BS detection and delay estimation are assumed to be complete, consistent with established generalized likelihood ratio test (GLRT) methods [4]. The scope is therefore narrowed to estimating the remaining parameters for the active set $\mathcal{K}$: CFOs $\boldsymbol{\omega}\triangleq[\omega_{k}]_{k\in\mathcal{K}}$, scaling factors $\boldsymbol{\mu}\triangleq[\mu_{k}]_{k\in\mathcal{K}}$, and per-frame channel gains $\boldsymbol{\alpha}\triangleq[\alpha_{k}^{(p)}]_{k\in\mathcal{K},\,p\in\mathcal{P}}$.
$\underset{\boldsymbol{\omega},\boldsymbol{\alpha},\boldsymbol{\tau},\boldsymbol{\mu}}{\text{maximize}}\quad\mathbb{P}\left(\mathbf{y}|\boldsymbol{\omega},\boldsymbol{\alpha},\boldsymbol{\mu}\right).$
@@
III Cross-Preamble Estimation
@@ III-A Channel Estimator
$\mathbf{y}^{(p)}=\underbrace{e^{j\omega(p-1)N_{\text{f}}}\boldsymbol{\Omega}\left(\boldsymbol{\omega}\right)\varodot\mathbf{C}\left(\boldsymbol{\tau}\right)}_{\boldsymbol{A}^{(p)}}\boldsymbol{\alpha}^{\left(p\right)}+\mathbf{v}^{(p)}$
where $\mathbf{C}\left(\boldsymbol{\tau}\right)=[\mathbf{0}_{\tau\times 1};\mathbf{c}_{k};\mathbf{0}]_{k\in\mathcal{K}}\in\mathbb{C}^{N_{\text{f}}\times K}$, $\boldsymbol{\Omega}\left(\boldsymbol{\omega}\right)=[\Omega_{k}(\omega_{k})]_{k\in\mathcal{K}}\in\mathbb{C}^{N_{\text{f}}\times K}$, $\Omega_{k}(\omega_{k})=[1;e^{j\omega_{k}};\cdots;e^{j\omega_{k}(N_{\text{f}}-1)}]$, $\boldsymbol{\alpha}^{(p)}=[\alpha_{k}^{(p)}]_{k\in\mathcal{K}}\in\mathbb{C}^{K\times 1}$, and $\mathbf{v}^{(p)}=[v[m]]_{m\in\{0,\cdots,N_{\text{f}}-1\}}\in\mathbb{C}^{N_{\text{f}}\times 1}$.
@@
$\mathbf{y}=\underbrace{\text{blkdiag}\left(\boldsymbol{A}^{(0)},\cdots,\boldsymbol{A}^{(P-1)}\right)}_{\boldsymbol{A}(\boldsymbol{\tau},\boldsymbol{\omega})}\boldsymbol{\alpha}+\boldsymbol{v}.$
Then, according to the LLS solution, the channel gain $\boldsymbol{\alpha}$ can be estimated by
$\hat{\boldsymbol{\alpha}}=\boldsymbol{A}(\boldsymbol{\tau},\boldsymbol{\omega})^{\dagger}\mathbf{y}.$ (5)
@@
$\boldsymbol{A}_{0}\boldsymbol{\alpha}+\underbrace{\left[\boldsymbol{A}^{(0)}\text{diag}\left(\boldsymbol{\alpha}^{\left(0\right)}\right);\cdots;\boldsymbol{A}^{(P-1)}\text{diag}\left(\boldsymbol{\alpha}^{\left(P-1\right)}\right)\right]}_{\boldsymbol{B}}\boldsymbol{\mu}+\boldsymbol{v}.$
where $\boldsymbol{A}_{i}=\text{blkdiag}(\boldsymbol{A}_{i}^{(0)},\cdots,\boldsymbol{A}_{i}^{(P-1)})$, $\boldsymbol{A}_{i}^{(p)}=e^{j\omega(p-1)N_{\text{f}}}$ $\boldsymbol{\Omega}(\boldsymbol{\omega})\varodot\mathbf{C}_{i}\left(\boldsymbol{\tau}\right)$, $\mathbf{C}_{i}\left(\boldsymbol{\tau}\right)=[\mathbf{0}_{(\tau+\tau_{\text{c}}i)\times 1};\mathbf{c}_{k,i};\mathbf{0}]_{k\in\mathcal{K}}$, and $\tau_{\text{c}}=N+\tau_{0}$. Then, according to the LLS solution, the scaling factor $\boldsymbol{\mu}$ can be estimated by
$\hat{\boldsymbol{\mu}}=\boldsymbol{B}^{\dagger}\left(\mathbf{y}-\boldsymbol{A}_{0}\right).$ (6)
III-B Separate Analysis and Separate CFO Estimator
@@
$r_{y,k}^{p,i}\left[\tau\right]\triangleq\frac{1}{N}\sum_{m=\tau}^{\tau+N-1}y\left[m+pN_{\text{f}}+i\tau_{\text{c}}\right]c_{k,i}^{\text{H}}\left[m-\tau\right]$ (7)
@@
$r_{y,k}^{p,i}\left[\tau_{k}\right]\sim\alpha_{k}^{(p)}r_{k}\left(\omega_{k}\right)\text{$\left(\mu_{k}e^{-j\omega_{k}\tau_{\text{c}}}\right)$}^{i}+\mathcal{CN}\left(0,\sigma_{k,p,i}^{2}\right)$ (8)
@@
Proof.
(Sketch) Substituting the received signal expression (4) into the cross-correlation definition (7), the resulting summation contains three parts:
i) Desired Signal: The auto-correlation yields the deterministic mean $\alpha_{k}^{(p)}r_{k}(\omega_{k})(\mu_{k}e^{-j\omega_{k}\tau_{\text{c}}})^{i}$.
ii) Interference: Cross-correlations with other transmitters $q\neq k$ produce a zero-mean complex Gaussian interference term with variance $\sum_{q\neq k,q\in\mathcal{K}}(\mu_{k}^{i}\alpha_{q}^{(p)}\sigma_{c})^{2}/N$.
iii) Noise: Correlation with noise contributes an additional zero-mean complex Gaussian noise term with variance $\sigma_{n}^{2}/N$.
Combining these contributions yields (8). ∎
@@
III-C Cross-Preamble CFO Estimator
Let the set of correlation variables associated with CFO $\omega_{k}$ as $\mathbf{r}_{y,k}=[r_{y,k}^{p,i}\left[\tau_{k}\right]]_{p\in\mathcal{P},i\in\{0,1\}}$, and define the cross-preamble likelihood as $\mathbb{P}(\mathbf{r}_{y,k}|\boldsymbol{\omega})$. Since preambles from different frames are non-overlapping, the correlation outputs are conditionally independent given $\omega_{k}$, yielding
$\mathbb{P}\left(\mathbf{r}_{y,k}|\boldsymbol{\omega}\right)=\prod_{p\in\mathcal{P}}\prod_{i\in\{0,1\}}\mathbb{P}\left(r_{y,k}^{p,i}\left[\tau_{k}\right]|\omega_{k}\right).$
@@ III-C Cross-Preamble CFO Estimator
$\hat{\omega}_{k}^{\text{cross}}\triangleq\arg\max_{\omega_{k}}\mathbb{P}\left(\mathbf{r}_{y,k}|\boldsymbol{\omega}\right)=\frac{1}{\tau_{c}}\angle\left\{\psi_{1}\psi_{0}^{\text{H}}\right\}$ (11)
@@
Proof.
(Sketch.) Since the autocorrelation function $r_{k}\left(\omega_{k}\right)$ depends on the CFO $\omega_{k}$, we introduce an auxiliary variable $\gamma_{k}\triangleq r_{k}\left(\omega_{k}\right)$ to absorb this dependence. Consequently, the unknown parameter vector is reparametrized as $\boldsymbol{\theta}\triangleq[\omega_{k},\gamma_{k},\gamma_{k}^{\text{H}}]^{\text{T}}$, where $\omega_{k}$ is real and $(\gamma_{k},\gamma_{k}^{\text{H}})$ is complex. Substituting into the log-likelihood function
$f\left(\boldsymbol{\theta}\right)$ $=\sum_{p\in\mathcal{P}}-\frac{\left|r_{y,k}\left[p,0\right]-\alpha_{k}^{(p)}\gamma_{k}\right|^{2}}{\sigma_{k,p,0}^{2}}$
$\quad+\sum_{p\in\mathcal{P}}-\frac{\left|r_{y,k}\left[p,1\right]-\alpha_{k}^{(p)}\gamma_{k}\mu_{k}e^{-j\omega_{k}\tau_{\text{c}}}\right|^{2}}{\sigma_{k,p,1}^{2}}.$
i) ML Estimator. Setting the gradients $\partial f/\partial\omega_{k}=0$ and $\partial f/\partial\gamma_{k}^{\text{H}}=0$ yields $\hat{\omega}_{k}^{\text{joint}}=\angle\left\{\psi_{1}\psi_{0}^{\text{H}}\right\}/\tau_{c}$.
ii) CRLB Analysis. The Fisher information matrix (FIM) is obtained from the second-order derivatives of $f\left(\boldsymbol{\theta}\right)$. Taking expectations over the noise and interference terms, the FIM is simplified to
$\mathbf{I}=\left[\begin{array}[]{ccc}2\tau_{\text{c}}^{2}\Gamma_{1}\left|\gamma_{k}\right|^{2}&j\tau_{\text{c}}\Gamma_{1}\gamma_{k}^{\text{H}}&-j\tau_{\text{c}}\Gamma_{1}\gamma_{k}\\ j\tau_{\text{c}}\Gamma_{1}\gamma_{k}^{\text{H}}&0&\Gamma_{0}+\Gamma_{1}\\ -j\tau_{\text{c}}\Gamma_{1}\gamma_{k}&\Gamma_{0}+\Gamma_{1}&0\end{array}\right].$
By applying the Schur complement, the CRLB for the CFO is given by
$\text{Var}\left(\hat{\omega}_{k}^{\text{joint}}\right)\geq\left[\mathbf{I}^{-1}\right]_{11}=\frac{\Gamma_{0}^{-1}+\Gamma_{1}^{-1}}{2\tau_{\text{c}}^{2}\left|r_{k}\left(\omega_{k}\right)\right|^{2}}$ (15)
@@
0:Initialization: $\hat{\boldsymbol{\omega}}\leftarrow\mathbf{0}$, $\hat{\boldsymbol{\alpha}}\leftarrow\mathbf{0}$, $\hat{\boldsymbol{\mu}}\leftarrow\mathbf{1}$, $\tilde{\mathbf{y}}\leftarrow\mathbf{y}$
1: Construct $\mathbf{A}$ based on $\hat{\boldsymbol{\omega}}$ and $\hat{\boldsymbol{\mu}}$ and update $\hat{\boldsymbol{\alpha}}$ based on (5); Then, Construct $\mathbf{A}_{0}$ and $\mathbf{B}$ based on $\hat{\boldsymbol{\omega}}$ and $\hat{\boldsymbol{\alpha}}$ and update $\hat{\boldsymbol{\mu}}$ based on (6).
2: For each BS $k$, perform SIC and compensation: $\tilde{\mathbf{y}}_{k}\leftarrow(\mathbf{y}-\boldsymbol{A}(\hat{\boldsymbol{\tau}}_{-k},\hat{\boldsymbol{\omega}}_{-k})\hat{\boldsymbol{\alpha}}_{-k})\varodot\Omega_{k}(\omega_{k})$. Then, estimate the remaining CFO $\delta_{k}$ based on the cross-preamble estimator (11), and update $\hat{\omega}_{k}\leftarrow\hat{\omega}_{k}+\delta_{k}$.
@@
III-D Joint CFO and Channel Estimation Algorithm
@@
$\tilde{\mathbf{y}}_{k}\;=\;\Bigl(\mathbf{y}-\boldsymbol{A}\bigl(\hat{\boldsymbol{\tau}}_{-k},\hat{\boldsymbol{\omega}}_{-k}\bigr)\hat{\boldsymbol{\alpha}}_{-k}\Bigr)\odot\boldsymbol{\Omega}_{k}(\hat{\omega}_{k})$ (16)
@@
Figure 2: CFO $\omega$ estimation NMAE over SINR, obtained from 200 Monte-Carlo simulations, with $K=12$ BSs, and $P=12$ preambles.
Figure 3: Channel $\alpha$ estimation NMSE over SINR, obtained from 200 Monte-Carlo simulations, with $K=12$ BSs, and $P=12$ preambles.
IV Simulation Results
@@
Fig. 2 illustrates the normalized mean absolute error (NMAE) over SINR
$\text{NMAE}=\frac{1}{f_{\text{scs}}}\frac{1}{N}\sum_{i=1}^{N}\frac{|\hat{\omega}_{i}-\omega_{i}|}{2\pi}$
where $f_{\text{scs}}$ is the subcarrier spacing (SCS).
As expected, error decreases with SINR. The ranking is consistent with the length of the used reference sequence: cross-preamble < separate < CP-based. Cross-correlation outperforms auto-correlation because the separation $\tau_{\text{c}}$ between the two training sequences provides a larger phase baseline. The proposed cross-preamble estimator further improves accuracy by pooling information across frames under the shared-CFO constraint. Our method achieves a normalized CFO error of approximately 0.05, even at -30 dB SINR. In OFDM systems, keeping the residual CFO well below 0.1 SCS is critical to minimize Inter-Carrier Interference (ICI). This result confirms that our estimator provides sufficient accuracy for reliable symbol demodulation. In contrast, the baseline (”Separate Analysis”) degrades significantly as the SINR drops, reaching a normalized error of 1.0 at -15 dB. An error of this magnitude implies a complete synchronization failure (e.g., integer subcarrier shift), rendering subsequent signal recovery impossible. Notably, the cross-preamble estimator at SINR $30\text{ dB}$ still outperforms the weighted-average baseline operating at SINR $5\text{ dB}$ confirming robustness in low-SINR regimes. Consequently, improving CFO accuracy reduces variance in channel estimation; the error (NMSE) versus SINR follows the same trend as in Fig. 3.
Figure 4: Aerial sampling campaign over the CUHK-Shenzhen campus. A DJI M300 RTK UAV equipped with an on-board Latte Panda 3 Delta computer, and a USRP B210 (calibrated) software-defined radio flies at an altitude of $150\text{ m}$ to collect measurements within the area outlined by the dashed red polygon. Red triangles with yellow labels indicate the locations and Cell ID of a subset of known terrestrial BSs.
Figure 5: Field test: Radio maps for beam 0 transmitted by a ground BS; the BS location is indicated by a red triangle.
Figure 6: Field test: Radio maps for beam 6 transmitted by a ground BS; the BS location is indicated by a red triangle.
We also verify our algorithm on the testbed, as shown in as illustrated in Fig. 4, a DJI M300 RTK UAV carries a universal software radio peripheral (USRP) to collect 5G NR signals, with carrier frequency $f_{c}=2.52495\text{ GHz}$ and bandwidth $20\text{ MHz}$.
Fig. 5 and 6 show the radio-map estimation for a low-altitude setting (2 of 8 beam settings). The two clearly resolved (pointing in different directions) beams corroborate the effectiveness of the proposed framework in practice. It is shown that the received power does not exhibit a simple monotonic distance- or propagation-loss trend but instead shows irregular spatial variations, since the dominant energy in the low-altitude airspace is largely contributed by ground-reflected paths.
V Conclusion